# Image-to-Video Transfer Learning based on Image-Language Foundation Models: A Comprehensive Survey

### v1

@@ V-D Video Question Answering
Pioneering methods [172, 173, 174, 175] either apply hierarchical spatio-temporal modeling strategies, employing two-stage architectures to first extract key frame clips followed by reasoning [176], or constructing spatio-temporal relational graphs via graph neural networks. Recently, more and more methods [120, 110] consider addressing the VideoQA problem utilizing the pre-trained foundation models.
@@
Many studies have focused on exploring frozen-feature strategies, such as knowledge distillation and post-network tuning, for image-to-video transfer learning. VideoDistill [34] introduces a language-aware gating module that dynamically selects and refines frame-level visual features. This design facilitates goal-oriented selective knowledge distillation and effectively mitigates language bias that often arises when adapting image-based models to video data. BIMBA [54] adopts the post-network tuning strategy by introducing a Mamba-based token compression module on top of pre-trained foundation models. The module models long-range spatiotemporal dependencies and compresses redundant tokens, thereby improving efficiency. LinVT[55] implements post-network tuning through a plug-and-play Linear Video Tokenizer module. Placed after a frozen visual encoder, this post-network employs a Spatio-Temporal Visual Token Refiner (SVR) and a Text-Conditioned Token Aggregator (TTA) to produce a compact, question-relevant video representation for the subsequent large language model. HierarQ [58] develops a post-network tuning module in the form of a trainable hierarchical Q-Former equipped with task-aware modulators and memory mechanisms. It generates a compact yet informative spatiotemporal representation, which is subsequently fed into a frozen large language model that is further adapted through a lightweight adapter for textual answer generation. TEA [86] extends TextVQA [177] from images to videos by introducing spatio-temporal recovery modules into a pretrained T5 language model. These modules, consisting of a temporal convolution block and an OCR-enhanced relative spatial bias, can capture spatio-temporal clues to support video question answering.
@@
Recent progress in Multimodal Large Language Models has established a new paradigm for Video QA, heavily reliant on the integration of modified visual and textual features. Atemporal Probe (ATP) [120] introduces a fine-tuning with prompt tuning approach that selects a single representative key frame to serve as a learnable visual prompt for a frozen vision encoder. Q-ViD [123] leverages a pre-trained InstructBLIP model [178] and incorporates tailored instructional prompts to generate descriptive video captions, using them as contextual cues for improved question answering. Similarly, ViTiS [122] employs multimodal prompt learning together with a visual mapping network to inject task-specific cues into image-language models, enabling strong performance with only few additional learnable parameters and limited training data. Other works explore adapter-based fine-tuning. Tem-Adapter [110] introduces a visual temporal aligner to predict future video states and a textual semantic aligner to refine embeddings for cross-modal alignment. TR-Adapter [118] proposes a temporal reasoning adapter designed to capture temporal relationships and facilitate knowledge transfer from language models. Unlike prior methods that maintain the original backbone architecture and perform adaptation through external lightweight modules, TC-LLaVA [84] modifies the attention computation in foundation models through partial tuning. Specifically, it introduces a Temporal-Aware Dual RoPE to encode temporal positional information and a Frame-wise Block Causal Attention Mask to enhance interactions among visual tokens.
@@
V-E General Video Understanding
Unlike most existing image-to-video transfer learning methods which are tailored for one specific task, there have been some methods that aim to transfer the pretrained image-text knowledge across multiple tasks to achieve general video understanding. FitCLIP [37] proposes to train a student model by a small number of labeled video-text pairs and a large number of video-text pseudo labels, where the student model is then fused with the teacher model by weight ensembling for zero-shot video understanding. STAN [64] presents spatial-temporal auxiliary network as a side-tuning approach to perform temporal interaction and feature merging in the additional side branch, which achieves state-of-the-art performances across video-text retrieval and video action recognition.
TABLE III: Experimental results of representative methods with varied image-language foundation models and different image-to-video transfer learning strategies in coarse-grained video understanding tasks. Coarse-grained Video Understanding Task Category Method Venue ILFM R@1 $\uparrow$ R@5 $\uparrow$ R@10 $\uparrow$ R@sum $\uparrow$ VTR Full Fine-Tuning CLIP4Clip [77] Neurocomp. 2022 CLIP 44.5 71.4 81.6 - Finetune with Prompt Tuning VOP [125] CVPR 2023 CLIP 44.6 69.9 80.3 - Finetune with Adapter MV-Adapter [168] CVPR 2024 CLIP 46.2 73.2 82.7 - Knowledge Distillation TeachCLIP [39] CVPR 2024 CLIP 46.8 74.3 - 203.7 Finetune with LoRA DiscoVLA [119] CVPR 2025 CLIP 50.5 75.6 83.8 209.9 Task Category Method Publication ILFM K400 Top-1 $\uparrow$ K400 Top-5 $\uparrow$ SSv2 Top-1 $\uparrow$ SSv2 Top-5 $\uparrow$ VAR Side-Tuning EVL [65] ECCV 2022 CLIP 87.3 - 66.7 - Finetune with Adapter AIM [113] ICLR 2023 CLIP 87.5 97.7 70.6 92.7 Finetune with Adapter DUALPATH [114] CVPR 2023 CLIP 87.7 97.8 71.4 93.4 Side-Tuning DiST [66] ICCV 2023 CLIP 88.0 97.9 73.1 93.2 Finetune with Adapter ZeroI2V [179] ECCV 2024 CLIP 87.2 97.6 72.2 93.0 Finetune with Adapter MoMa [102] ICML 2025 CLIP 87.8 98.0 73.8 93.6 Task Category Method Publication ILFM MSRVTT-MC $\uparrow$ NextQA $\uparrow$ MSVD-QA $\uparrow$ How2QA $\uparrow$ VideoQA Finetune with prompt tuning ATP[120] CVPR 2022 CLIP 93.2 54.3 - 65.1 Finetune with Adapter Tem-Adapter[110] ICCV 2023 CLIP 94.3 - - - Finetune with prompt tuning ViTiS[122] ICCV 2023 CLIP - - 47.8 - Finetune with prompt tuning Q-Vid[123] ACL 2024 CLIP - 66.3 - 71.4 Knowledge Distillation VideoDistill[34] CVPR 2024 CLIP 97.8 - 49.2 - Post Network Tuning BIMBA[54] CVPR 2025 LLaVA - 83.7 - - Task Category Method Publication ILFM BLEU4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$ Video Captioning Post Network Tuning CLIP4Caption[56] ACM MM 2021 CLIP 47.2 64.8 31.2 60.0 Full Fine-Tuning SWINBERT[69] CVPR 2022 CLIP 45.4 64.1 30.6 55.9 Finetune with partial tuning mPLUG-2[87] ICML 2023 CLIP 57.8 70.1 34.9 80.3 Fine-Tuning with Extra Models Vid2Seq[103] CVPR 2023 CLIP - - 30.8 64.6 Side-Tuning MAMS[67] AAAI 2025 - 60.0 71.2 34.7 82.9 Post Network Tuning HierarQ[58] CVPR 2025 LLaVA - - 35.1 80.5
VI Experimental Analysis
@@
VI-A Temporal Video Grounding
Datasets and Metrics. Temporal Video Grounding (TVG) aims to locate segments (or moment) in given untrimmed video based on a descriptive sentence. To address the TVG problem, numerous datasets have been constructed with video–text pairs and detailed annotations specifying the temporal spans of events described in the text. Specifically, DiDeMo[149], TACoS[180], Charades-STA[148] and ActivityNet Captions[181] are three most widely used benchmarks in the community. Following [148], most existing approaches are evaluated in these datasets and report their performance in the term of metrics ”Rn@m”and ”mIoU”. Metric ”Rn@m” is employed to compute the the percentage of samples where the top-$n$ predicted segments has an IoU greater than a predefined threshold $m$, where $m$ is typically defined as $\{0.1,0.3,0.5,0.7\}$. Metric ”mIoU” calculates the mean IoU between the predicted segments and ground truth on all testing data.
Results and Analysis. As shown in Table III, results for Temporal Video Grounding (TVG) reveal a notable performance gap between specialized architectures built on CLIP and generalist Multimodal Large Language Models (MLLMs) adapted from LLaVA. Among CLIP-based models, the Side-Tuning method R2-tuning achieves the highest precision with 59.8% on R@0.5 and 50.9% on m_I0U. This can be attributed to its ability to refine multi-level visual features within a dedicated temporal modeling module. In contrast, for MLLM-based methods, NumPro leads with 42.0% on R@0.5 and 41.4% on m_IoU by employing a Fine-Tuning with Extra Models strategy that reformulates temporal regression into a visual OCR task, better aligning with the MLLM’s native strengths. The observed results suggest that for TVG—a task requiring precise temporal regression—specialized, decoupled architectures currently hold an advantage over temporally-augmented general-purpose MLLMs.
VI-B Video-Text Retrieval
Datasets and Metrics. Video Text Retrieval (VTR) aims to identify relevant videos from a large collection based on a given descriptive sentence, or vice versa. To facilitate research in this area, several datasets have been constructed, pairing videos with textual descriptions and enabling the evaluation of cross-modal alignment performance. Among these, MSR-VTT [182] is one of the most widely used benchmarks in the field. Following common practice, most existing approaches are evaluated on this dataset and report performance using the metrics R@1, R@5, R@10, and R@sum. These metrics are defined as follows: R1, R5, and R10 measure the recall rate at the top 1, 5, and 10 retrieved results, respectively, indicating the model’s ability to rank relevant items at the top. Rsum is the sum of R1, R5, and R10, providing an aggregated measure of retrieval performance across different ranks.
Results and analysis. By examining the results presented in Table II for the task of Video-Text Retrieval, it can be observed that most recent methods are built upon the pretrained CLIP model due to its strong visual-textual correspondence capability. Among these, fine-tuning with LoRA achieves the best performance, significantly surpassing other feature-modification transfer learning strategies such as prompt tuning, knowledge distillation, and adapter-based tuning, in the metric of R@1. As expected, directly fusing CLIP-extracted features with a post-network tuning scheme performs worse than most of the approaches that updates CLIP parameters, highlighting the importance of adapting the original foundation model to the target task and dataset for optimal performance.
VI-C Video Action Recognition
Datasets and Metrics. Video Action Recognition (VAR) is a classical task in video understanding, which aims to recognize the category of human action being performed in a trimmed video. To support research in this area, several influential datasets have been developed. Foundational benchmarks include UCF101 [183] and HMDB51 [184], which established early standards for the task. More recently, the large-scale Kinetics series(e.g., Kinetics-400 [185], Kinetics-600 [186], and Kinetics-700 [187]) has become the most widely adopted benchmark for evaluating modern, high-capacity models. Following common practice, model performance on these datasets is primarily evaluated using Top-1 and Top-5 accuracy. Top-1 accuracy measures the percentage of videos for which the model’s single highest-probability prediction is correct. Top-5 accuracy measures the percentage of videos where the ground-truth label is among the model’s five most confident predictions, which is particularly informative for datasets with a large number of classes.
Results and analysis. Table II reports a detailed comparison of several representative image-to-video transfer methods for video action recognition based on the CLIP model. Overall, finetuning with Adapter strategy achieves better results than side-tuning strategy in this task. MoMa [102] obtains the best results on both K400 and SSv2 sets, which employs a mamba architecture to learn long-term fine-grained spatio-temporal information. The comparison between EVL and DiST demonstrates the effectiveness of disentangling the spatial and temporal modeling in image-to-video transfer learning.
VI-D Video Question Answering
Datasets and Metrics. Video Question Answering (VideoQA) task is typically evaluated using metrics such as Top-1 Accuracy, Top-K Accuracy, and Average Normalized Levenshtein Similarity (ANLS)[188], which measure the alignment between the predicted answers and ground truth answers. The VideoQA task is often subdivided into three distinct subcategories based on the nature of the questions and the expected answers: i):Multiple-Choice (MC): This subcategory involves answering questions that have a predefined set of possible answers. Common benchmarks for evaluating MC-based VideoQA tasks include TGIF-Action and TGIF-Transition[172], which focus on action-based queries within video clips, and MSRVTT-MC [182] and LSMDC-MC [189], which extend this challenge to larger, more diverse video datasets. ii):Open-Ended (OE): In this subcategory, the questions posed are more open-ended, allowing for subjective, creative, and sometimes logical responses. Popular benchmarks in the OE VideoQA subtask include TGIF-Frame [172], which deals with frame-level visual understanding; MSRVTT-QA [182], which presents questions based on the MSR-VTT video dataset; MSVD-QA [190], which uses the MSVD dataset for question answering; and LSMDC-FiB [189] and ActivityNet-QA [191], which extend the challenge to more diverse and dynamic video sequences. iii):Long-Form (LF): The Long-Form VideoQA[192] task takes VideoQA a step further by not only requiring a single answer but also necessitating the generation of comprehensive, detailed explanations. Key benchmarks in this subtask include ActivityNet-QA [191], How2QA [193], and VIOLIN [194], all of which emphasize the generation of detailed, explanatory answers that require a deep understanding of video context and reasoning across time.
Results and analysis. We present the comparison results on four datasets in Table II, where most competitors use CLIP as their backbone to ensure a fair comparison. For standard vision-language models such as CLIP, various transfer methods demonstrate strong and dataset-dependent performance. The distillation-based approach, VideoDistill [34], achieves state-of-the-art results on MSRVTT-MC and MSVD-QA with accuracies of 97.8% and 49.2%, respectively. In contrast, on the How2QA benchmark, the prompt-tuning method Q-Vid [123] attains the highest score at 71.4%. These results suggest that for a fixed backbone like CLIP, the most effective adaptation strategy varies across different benchmarks.
Conversely,the integration of Large Language Models as the foundation model substantially elevates performance. On the NextQA benchmark, the LLaVA-based BIMBA [54] model, which utilizes a post-network tuning strategy, achieves an accuracy of 83.73%. This represents a significant margin of improvement over the best CLIP-based method, Q-Vid with an accuracy of 66.3% on the same dataset. This finding underscores that post-network tuning is a potent transfer strategy for LLM-based models, as it efficiently adapts visual information to unlock the LLM’s powerful, built-in semantic reasoning for video understanding.
VI-E Video Captioning
Datasets and Metrics. A variety of standard benchmarks are collected to train and evaluate video captioning models. Prominent open-domain datasets include MSRVTT [195], comprising 10,000 video clips across 20 categories, and MSVD dataset [196]. Both datasets provide multiple human-annotated reference captions per video, making them ideal for evaluating models’ general-purpose descriptive capabilities and serving as foundational benchmarks for assessing performance across diverse scenarios. For caption tasks that require specialized procedural understanding, datasets such as YouCook2 [197] are commonly used. This dataset contains long instructional videos with temporally segmented annotations, challenging models to generate fine-grained, step-by-step descriptions of procedural content.
The performance of video captioning models is typically evaluated by comparing machine-generated candidate captions with a set of human-written reference captions using a suite of automated metrics. These metrics primarily measure both lexical similarity and semantic alignment between the generated and reference captions. For instance, BLEU@N [198] is a precision-oriented metric that measures n-gram overlap to evaluate fluency, while the recall-oriented ROUGE-L [199] uses the Longest Common Subsequence (LCS) to assess content fidelity. Building on these, METEOR [200] achieves better agreement with human judgment by considering synonyms and word stems, providing a more semantically informed evaluation of generated captions. CIDEr [201], developed specifically for captioning, measures consensus by representing sentences as TF-IDF–weighted n-gram vectors, rewarding captions that capture the salient information shared across the reference set. These metrics provide a holistic evaluation of caption quality, capturing both linguistic accuracy and semantic fidelity.
Results and analysis. As shown in Table II, recent advancements in video captioning are driven by modified transfer strategies that outperform traditional fine-tuning. The side-tuning method MAMS [67] achieves state-of-the-art performance across multiple key metrics (60.0 BLEU4, 71.2 ROUGE-L, 82.9 CIDEr). This success can be primarily attributed to its architectural design; by training a separate, lightweight module while keeping the foundation model frozen, it effectively learns task-specific temporal dynamics without the risk of catastrophic forgetting. The strong performance of the LLaVA-based HierarQ [58], which excels by aligning visual features with an LLM’s powerful reasoning, further validates this trend. These results indicate that the most effective transfer strategies are those that intelligently integrate new knowledge while preserving the core capabilities of the pre-trained backbone.
VI-F Spatio-Temporal Video Grounding
Datasets and Metrics. For the STVG task, VidSTG is currently the most widely-used benchmark dataset. It contains 99,943 video-text pairs, including 44,808 declarative sentences and 55,135 interrogative sentences from 10,303 videos spanning 80 object categories. For clearer comparison, we evaluate different methods under the declarative sentences setting. As for metrics, we follow previous works[91, 92] to use mean vIoU as the main metric. vIoU is computed as: $vIoU=\frac{1}{\left|T_{u}\right|}\sum_{t\in T_{i}}IoU(\hat{b}_{t},b_{t}),$ where $T_{i}$ and $T_{u}$ indicate the intersection and union between the time intervals obtained from annotation and prediction, respectively. $\hat{b}_{t},b_{t}$ are the predicted and ground truth bounding box for the $t$-th frame, respectively. We average the vIoU score over all the samples to obtain mean vIoU (termed m_vIoU). We also report vIoU@R which indicates the proportion of samples with vIoU higher than $R$. We also follow TubeDETR[70] and report the tIoU evaluating the temporal grounding performances. It is defined as: $tIoU=\frac{\left|T_{i}\right|}{\left|T_{u}\right|},$ where $T_{gt}$ indicates the set of frames inside the ground truth time interval.
Results and analysis. In fully-supervised STVG tasks, nearly all existing approaches adopt MDETR[163] as the foundation model, as it provides essential visual grounding capabilities and can be fully adapted to video-level understanding through fine-tuning. However, as shown in Table III, performing temporal modeling solely on the image-level features extracted by MDETR leads to limited temporal localization accuracy. Consequently, recent methods [89, 90, 91, 92] have incorporated auxiliary models to extract video-level features, such as SlowFast[137] and 3D-CNN[164], which substantially improve temporal localization performance, as evidenced by the increase in tIoU from 48.1 in TubeDETR [70] to 52.1 in CoSTA [90]. These results suggest that integrating both image-level and video-level features enables the model to achieve more precise fine-grained spatio-temporal localization. However, incorporating auxiliary models also increases computational cost; for example, TASTVG[92] was trained on 32 NVIDIA A100 GPUs for 15 hours.
VI-G Open Vocabulary Video Instance Segmentation
Datasets and Metrics. We compare the experimental results of all OVVIS methods on the LV-VIS[43] validation and test sets. LV-VIS is a large-vocabulary video instance segmentation dataset comprising 4,828 real-world videos spanning 1,196 categories. The videos are partitioned into 3,083 for training, 837 for validation, and 908 for testing. The categories are divided into 641 base categories—inherited from frequent and common classes in LVIS[202]—and 555 novel categories that are disjoint from the base categories. We adopt the Average Precision (AP) as the evaluation metric, computed as the mean over multiple IoU thresholds from 0.5 to 0.95. To better assess open-vocabulary performance, we also report APn, which denotes the average AP exclusively on novel categories.
Results and analysis. As shown in Table III, Recent OVVIS methods have adopted the paradigm of post-network tuning for image-to-video transfer learning, incorporating CLIP to introduce open-vocabulary capabilities. Through a comparative analysis of several CLIP-based methods, we observe that CLIP-VIS[44] achieves the best performances. This can be primarily attributed to its introduction of a temporal Top-K enhanced matching strategy. Instead of matching only adjacent frames or all historical frames, this approach selects the top K most similar frames from the past for query association. This mechanism reduces interference from noisy frames and mitigates error propagation, thereby enhancing tracking stability. Furthermore, the method integrates classification scores with an objectness score and a mask IoU score. This fusion strengthens the consistency between classification and segmentation quality, improving robustness against challenges such as occlusion and complex scenes.
VI-H Open Vocabulary Multi-Object Tracking
Datasets and Metrics. For the OVMOT task, we adopt the TAO[203] dataset as benchmark and compare the performance of existing image-to-video transfer methods. The TAO dataset is a comprehensive video collection designed to provide rich visual and semantic annotations for diverse video analysis tasks. A key characteristic of TAO is its open-world nature, enabling the tracking of arbitrary objects rather than restricting to predefined categories. The dataset is partitioned into three subsets: training, validation, and testing, comprising 500, 988, and 1,419 videos respectively.
The OVMOT method is primarily evaluated using the TETA metric. TETA[204], an extension of HOTA[205], enhances its functionality to address challenges in long-tail scenarios. The TETA score comprises three components: a localization score (LocA), a classification score (ClsA), and an association score (AssocA). LocA quantifies the precision of bounding box localization, measuring the spatial alignment between predicted and ground-truth boxes, irrespective of classification or association accuracy. ClsA evaluates the correctness of object classification into predefined categories (e.g., ’car’, ’pedestrian’, ’bicycle’), independent of localization or association performance. AssocA assesses the robustness of identity preservation over time, reflecting the consistency of object-ID assignments across frames and directly capturing tracking and data association efficacy. The overall TETA score is computed as the arithmetic mean of these three sub-scores.

### v2

@@ V-D Video Question Answering
Pioneering methods [69, 43, 27, 187] either apply hierarchical spatio-temporal modeling strategies, employing two-stage architectures to first extract key frame clips and then perform reasoning [82], or construct spatio-temporal relational graphs using graph neural networks. Recently, more and more methods [9, 16] consider addressing the VideoQA problem utilizing the pre-trained foundation models.
@@
Many studies have focused on exploring frozen-feature strategies, such as knowledge distillation and post-network tuning, for image-to-video transfer learning. VideoDistill [219] introduces a language-aware gating module that dynamically selects and refines frame-level visual features. This design facilitates goal-oriented selective knowledge distillation and effectively mitigates language bias that often arises when adapting image-based models to video data. BIMBA [68] adopts the post-network tuning strategy by introducing a Mamba-based token compression module on top of pre-trained foundation models. The module models long-range spatiotemporal dependencies and compresses redundant tokens, thereby improving efficiency. LinVT [45] implements post-network tuning through a plug-and-play Linear Video Tokenizer module. Placed after a frozen visual encoder, this post-network employs a Spatio-Temporal Visual Token Refiner (SVR) and a Text-Conditioned Token Aggregator (TTA) to produce a compact, question-relevant video representation for the subsequent large language model. HierarQ [2] develops a post-network tuning module in the form of a trainable hierarchical Q-Former equipped with task-aware modulators and memory mechanisms. It generates a compact yet informative spatiotemporal representation, which is subsequently fed into a frozen large language model that is further adapted through a lightweight adapter for textual answer generation. TEA [208] extends TextVQA [153] from image to video by incorporating a spatio-temporal recovery module, which includes a temporal convolution block and an OCR-enhanced relative spatial bias, into pretrained T5 language model.
@@
Recent progress in MLLMs has established a new paradigm for Video QA, heavily reliant on the integration of adapted visual and textual features. Atemporal Probe (ATP) [9] introduces a fine-tuning with prompt tuning approach that selects a single representative key frame to serve as a learnable visual prompt for a frozen vision encoder. Q-ViD [133] leverages a pre-trained InstructBLIP model [28] and incorporates tailored instructional prompts to generate descriptive video captions, using them as contextual cues for improved question answering. Similarly, ViTiS [37] employs multimodal prompt learning together with a visual mapping network to inject task-specific cues into image-language models, enabling strong performance with only few additional learnable parameters and limited training data. Other works explore adapter-based fine-tuning. Tem-Adapter [16] introduces a visual temporal aligner to predict future video states and a textual semantic aligner to refine embeddings for cross-modal alignment. TR-Adapter [180] proposes a temporal reasoning adapter designed to capture temporal relationships and facilitate knowledge transfer from language models. Unlike prior methods that maintain the original backbone architecture and perform adaptation through external lightweight modules, TC-LLaVA [46] modifies the attention computation in foundation models through partial tuning. Specifically, it introduces a Temporal-Aware Dual RoPE to encode temporal positional information and a Frame-wise Block Causal Attention Mask to enhance interactions among visual tokens.
VI General Video Understanding
Unlike most existing image-to-video transfer learning methods, which are tailored for one specific task, there have been some methods that aim to transfer the pretrained image-text knowledge across multiple tasks to achieve general video understanding as a video-language foundation model.
FitCLIP [13] proposes to train a student model by a small number of labeled video-text pairs and a large number of video-text pseudo labels, where the student model is then fused with the teacher model by weight ensembling for zero-shot video understanding. STAN [115] presents a spatial-temporal auxiliary network as a side-tuning approach to perform temporal interaction and feature merging in the additional side branch, which achieves SOTA performances across video-text retrieval and video action recognition. Video-ChatGPT [130] employs a frozen CLIP-L/14 visual encoder to extract both spatial and temporal video features by averaging frame-level representations across their respective dimensions. The resulting spatiotemporal features are subsequently projected into the large language model’s input space via a learnable linear layer through video instruction tuning. Video-LLaMA [203] incorporates an additional Video Q-Former with temporal position encoding and a frozen LLM as the post-network following a pretrained frozen BLIP-2 [89] visual encoder. The architecture also introduces a similarly structured audio branch, enabling the model to process both visual and auditory modalities simultaneously. Video-LLaVA [106] utilizes LanguageBind [215], initializing from Open-CLIP as a frozen video encoder, and subsequently fine-tunes the LLM and projection layer following the methodology established in LLaVA[114]. UniformV2 [91] demonstrates that incorporating specifically designed local and global relation aggregators after a large-scale image-text pre-trained visual encoder, followed by video pre-training, can significantly enhance performance. Based on CLIP’s multimodal contrastive learning with video data, InternVideo [178] incorporates an additional masked video encoder. This encoder learns meaningful spatiotemporal representations by reconstructing masked video tokens. Furthermore, the framework unifies their representations through lightweight interactive model supervision. Experiments confirm that features learned via generative and contrastive training are complementary, and their integration yields superior performance compared to training either objective independently. InternVideo2 [177] employs InternVL [25] as the teacher model, utilizing a distillation paradigm to transfer knowledge from the ILFMs to the video encoder before proceeding with subsequent multimodal alignment and LLM-linking. InternVideo2.5 [179] employs InternVL2.5 [24] as its visual encoder and introduces hierarchical token compression alongside task preference optimization during pre-training, enabling robust long and rich context understanding in MLLMs.
@@ VI General Video Understanding
VII Experimental Analysis
@@
VII-A Open Vocabulary Multi-Object Tracking
Datasets and Metrics. For the OV-MOT task, we use TAO[29] as the benchmark to compare existing image-to-video transfer methods. TAO is an open-world video dataset with detailed annotations, supporting tracking of arbitrary objects. This set is divided into three subsets: training, validation, and testing, comprising 500, 988, and 1,419 videos, respectively. The OV-MOT method is primarily evaluated using the TETA metric. TETA[96], an extension of HOTA[124], desiged for long-tail scenarios. TETA comprises three components: the localization score (LocA), which measures bounding box alignment; the classification score (ClsA), which evaluates object category correctness; and the association score (AssocA), which assesses ID consistency across frames. The overall TETA score is the arithmetic mean of these three sub-scores.